ON COMBINATORIAL CALCULATIONS FOR THE JAMES–HOPF MAPS

Jie Wu

Abstract. We give some formulas of the James-Hopf maps by using combinatorial methods. An application is to give a product decomposition of the spaces ΩΣ^{2}(X).

Typeset by AMS-TEX